Shift Summary for day shift, 24/06/99

Ashkan_Alavi-Harati (alavi@fnal.gov)
Thu, 24 Jun 1999 15:31:19 -0500

Shift Summary for day shift, 24/06/99

On Shift: Brad Cox, Ashkan Alavi
Also present: All the Arizona group, Ed, Hogan, Marj, Patty,
and others.

Hours of beam delivered at all intensities: 0
Hours of beam delivered at nominal intensity: 0
Hours of E832 data (special runs not included): 0
Hours of special runs (laser scans, muon runs...): 3

Accounting of lost time:

Tevatron scheduled down time for the ground fault repair.

========================================================================
Runs ended this shift:
Run 13706 - 16.0M events - BEAM_832
Run 13709 - Laser_Scan
Run 13715 - KTEV_PED
Run 13716 - KTEV_PED

Tapes brought to FCC this shift:
QKC 37 38 39
QKS 37 38 39
QKK 37 38 39
QKB 13

========================================================================
Accesses during shift: 3
Marj and her student Colin to swap DC1 wires.
Valeri and Ashkan to adjust VV' high voltage.
Ed and Ashkan to swap four DPMT's.
Marj and Angela to sample DC gas.
ED, Bob T and Konrad K.(from NA48) for a quick tour!
========================================================================
Overview of shift:

Making access to fix DC1 wires and swap bad DPMTs and do other repair
works during the scheduled down time.
E832 data taking, if we get any beam.

========================================================================
Description of Shift:

run spill time comment
------------------------------------------------------------------------
13706 481 07:01 Beam has gone away as scheduled.
13706 485 07:25 Stop the run. The run didn't stop properly
(hanging for about 20 min), saved the histos,
did a fresh start and installed fresh tapes.
13715 13:40 Run starts.
13715 14:20 Run ends.
13716 14:30 Run starts.
13716 14:20 Run ends.

Others:

10:05 EB and ECC removed 1 nsec of delay from RF.

10:19 ECC and Jianbo swapped an HCC board and fixed the HCC error.

10:35 VJ raised the high voltage of 9 VV' channels by 20 V.

11:30 MDC and CW swapped DC1YD wires 5 and 6 back to their proper
positions. They also measured the z position of the Regenerator, DC1 and
DC2 as follows:
REG: 3.9930
DC1: 38.5450
DC2: 44.6621

12:00 EB swapped four DPMTs: 1642, 1678, 2310, 2589

13:40 ST installed v313 of Run Control which correctly writes tape labels
without DRT_FRESH_START.

13:50 MDC and AB took gas samples from all drift chambers for GC
analysis.

14:25 Ed installed new CSQ6 constants based on run 13698, and also
modified channel 1642. New constants will be valid for runs from 13716.
Furthermore, new peds installed for CsI ADCs from run 13715.

14:45 A visit by the Computing division operating system support group to
see where the online machines are, in case the new personnel gets page in
the middle of the night!

15:30 Still no beam. Estimated to get beam within next hour.

========================================================================
Follow up Notes on Previous Shift

========================================================================
Items to Follow up on Next Shift